OpenStudy (anonymous):

describe the set C f: I ---> R for which \[\forall x,y \in I \left| f(x)-f(y)\right| \le \sin ^{2} (x-y) \] please help !

OpenStudy (anonymous):

\(\large C:= \lbrace f: I \to \mathbb{R}, \ \forall x,y \in I \mid |f(x)-f(y)| \leq \sin^2(x-y) \rbrace \) ?

OpenStudy (anonymous):

yes it is.. do you have an idea ?

OpenStudy (anonymous):

I am not sure what they want you to do when they say describe the set, check if the set is bound/unbound, open/closed, empty/nonempty?

OpenStudy (anonymous):

any idea ?